OpenStudy (anonymous):

how do i work out which number is greater? (sqrt12 - sqrt11), (sqrt13 - sqrt12)

OpenStudy (asnaseer):

@tomiko are you aware of this factorisation?\[a^2-b^2=(a+b)(a-b)\]

OpenStudy (anonymous):

yes. but i don't know how to apply this to work out the answer. please help me

OpenStudy (asnaseer):

look at the first term:\[\sqrt{12}-\sqrt{11}\]if you multiply this by \(\sqrt{12}+\sqrt{11}\) what will you get?

OpenStudy (anonymous):

you get 1

OpenStudy (asnaseer):

correct. nor similarly multiply the second expression with \(\sqrt{13}+\sqrt{12}\) - what do you get?

OpenStudy (asnaseer):

*now

OpenStudy (anonymous):

for that too i get 1. but one is definately greater. cause when i use the calculate i get 2 different values, making one of the pairs greater than the other.

OpenStudy (asnaseer):

good. now, out these two terms, which one is bigger:\[1. \sqrt{12}+\sqrt{11}\]\[2.\sqrt{12}+\sqrt{13}\]

OpenStudy (anonymous):

\[\sqrt{12} + \sqrt{13}\] is bigger

OpenStudy (asnaseer):

good. so you now know that: a) x times (1) = 1 b) y times (2) = 1 and you know (2) is bigger than one. therefore what can you conclude about x and y?

OpenStudy (asnaseer):

I meant to say: you know (2) is bigger than (1)

OpenStudy (anonymous):

yes. 2 is bigger than one so that makes \[\sqrt{13} - \sqrt{12}\] bigger?? :D

OpenStudy (asnaseer):

no - in terms of x and y that I listed, x is being multiplied by a smaller number than y is being multiplied by, but both end up as 1. therefore is x or y bigger?

OpenStudy (anonymous):

in that case x is bigger

OpenStudy (asnaseer):

imagine you had:\[x*4=1\]\[y*8=1\]which one would you say is bigger, x or y?

OpenStudy (anonymous):

x

OpenStudy (asnaseer):

yes, x is bigger, therefore you know which one is bigger in your original question

OpenStudy (anonymous):

yes i know \[\sqrt{13} - \sqrt{12}\] is bigger.

OpenStudy (asnaseer):

look back carefully through what was derived and re-think your answer.

OpenStudy (asnaseer):

remember that \(\sqrt{13} - \sqrt{12}\) was multiplied by the bigger number \(\sqrt{13} + \sqrt{12}\) to get 1.

OpenStudy (anonymous):

yes!! i get it now. had to look at my solutions again using the different of 2 squares.

OpenStudy (asnaseer):

we showed that:\[(\sqrt{12}-\sqrt{11})(\sqrt{12}+\sqrt{11})=1\]\[(\sqrt{13}-\sqrt{12})(\sqrt{13}+\sqrt{12})=1\]and\[\sqrt{13}+\sqrt{12}\gt\sqrt{12}+\sqrt{11}\]

OpenStudy (anonymous):

that makes \[\sqrt{12} - \sqrt{11}\] greater!

OpenStudy (asnaseer):

perfect! well done! :)

OpenStudy (anonymous):

i undertand now. thank you very much! but you have any idea about this approach? Purely algebra ? Place ? i/o =, > or <. sqrt(12)-sqrt(11) ? sqrt(13)-sqrt(12) 2sqrt(12) ? sqrt(11)+sqrt(13) 48 ? ... Can you continue ?

OpenStudy (asnaseer):

ok what this approach is doing is bringing together the two square roots of 12 and and then squaring both sides. so:\[(2\sqrt{12})^2=4*12=48\]now you just need to square the right hand side:\[(\sqrt{11}+\sqrt{13})^2=?\]

OpenStudy (asnaseer):

do you understand?

OpenStudy (anonymous):

yeap...! i'm trying to do it on paper here :D

OpenStudy (asnaseer):

ok :)

OpenStudy (anonymous):

and this shows \[\sqrt{12} - \sqrt{11}\] is bigger. you can see my working attached as an image

OpenStudy (asnaseer):

you expansion of \((\sqrt{11}+\sqrt{13})^2\) is not correct. remember that:\[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

so if i correct it. i get \[48 ? 24 + 2\times \sqrt{13}\times \sqrt{11}\]

OpenStudy (asnaseer):

that is correct so far. now take the 24 to the left hand side and then divide both sides by 2

OpenStudy (asnaseer):

also remember that:\[\sqrt{a}\times\sqrt{b}=\sqrt{ab}\]

OpenStudy (anonymous):

now i get (attached)

OpenStudy (asnaseer):

that is correct - just one step left now - square both sides BTW: It is better to save your images as .png format if you can - they are smaller than .jpg files

OpenStudy (anonymous):

now the final answer!! :d

OpenStudy (asnaseer):

great! I'm glad you finally got there :)

OpenStudy (anonymous):

attached

OpenStudy (asnaseer):

perfect!

OpenStudy (anonymous):

sorry for all the bother. thanks for the help. i finished high school 6 years ago and didn't do any math at all and now just started uni. it's going to be a little change for me. i'm doing computer science now. lots of math in my class!! but i love subject.

OpenStudy (asnaseer):

no problem at all my friend - the main thing here is you obviously love to learn! :)

OpenStudy (anonymous):

challenge** thank you very much once again. now any question like this that comes my way i will murder it!!

OpenStudy (asnaseer):

he he - I love your "ambition" - keep up the good work! :)